OpenStudy (anonymous):

Solve for b: 5b^2-5b+1=0

OpenStudy (saifoo.khan):

Use the Quadratic Equation.

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=5,b=-5,c=1\]

OpenStudy (anonymous):

I used that and the answer came out to be x=+/-square root 5/5. Does this sound crazy

OpenStudy (saifoo.khan):

No it dosn't.

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

Actually when you use the quadratic formula the answer you should get is: \[\ \frac{5\pm \sqrt{5}}{10} \]

OpenStudy (anonymous):

Ok. So I don't need to reduce it any further?

OpenStudy (anonymous):

no, you only reduce it further if all of the numbers there share a common factor say for instance if it was \[\ \frac{5\pm5\sqrt{5}}{10} \]

OpenStudy (anonymous):

You know you have a common factor 5 that you can reduce

OpenStudy (anonymous):

ok. THx

OpenStudy (anonymous):

you're welcome